A political candidate has asked you to conduct a poll to determine what percentage of people support her.

If the candidate only wants a 1% margin of error at a 90% confidence level, what size of sample is needed?

Give your answer in whole people
Answer #1

Solution :

Given that,

\hat p = 0.5

1 - \hat p = 0.5

margin of error = E = 0.01

Z\alpha/2 = 1.645

sample size = n = (Z\alpha / 2 / E)2 * \hat p * (1 - \hat p )

= (1.645 / 0.01)2 * 0.5 * 0.5

= 6766

sample size = n = 6766